# After the Breach: Incident Response within Enterprises

### v1

@@ 1. Introduction
Once inside, attackers use a variety of techniques to move laterally through the network (e.g., pass-the-hash, Kerberoasting, etc), compromise accounts with high privileges and access to sensitive systems or data, and use those privileges to ferry data outside the network. Throughout this process, attackers play a carefully crafted cat-and-mouse game against threat detection systems, trying to masquerade as legitimate users within the network (e.g., by using legitimate tools, or testing access to systems infrequently). Indeed, this process can often be long, and as a result, attackers can persist inside the network for weeks, if not months before being detected. [^2] In fact, IBM reports that the average time to detect and contain a data breach in 2023 is 204 days (IBM, 2023) For example, the 2020 SolarWinds hack persisted in the network for 14 months before being detected. The 2023 HHS ransomware attack persisted in the network for 9 months before being detected. The 2023 23andMe breach took place over a duration of 5 months before being detected.
@@
Enterprises too employ various systems to detect and respond to such attacks – those that monitor endpoint or host activities (e.g., EDS: Endpoint Detection and Response systems), those that monitor network traffic (e.g., Intrusion and Detection Systems: NIDS, or Intrusion Prevension Systems: IPS), monitor user activity (e.g., User and Entity Behavior Analytics: UEBA), aggregate and monitor system information such as logs (e.g., Security Information and Event Management - SIEM). These systems are often deployed on endpoint devices (e.g., servers, laptops), inside network infrastructures (e.g., routers, switches), or in centralized hardware (e.g., Active Directory server), and are passively monitoring for threats. These systems are deployed in a hierarchical manner, in conjunction with each other, with a team of analysis to triage and respond to threats. It is this step that is labor intensive and time consuming, and often has the highest risk of missing a true positive (commonly termed the needle-in-the-haystack problem). As a result, several systems have been proposed to aid the investigative process by tracking related events, piecing together a causal reasoning of those events that led to the attack (termed ”causal analysis”), and understanding the scope of the attack, all in an effort to minimize false positives, and accelerate the triage process.
@@
Unfortunately, the scale and complexity of an enterprise poses several challenges in building such systems. First, the sheer volume of data that needs to be analyzed is large. Second, the data is often noisy, and is heavily skewed towards benign activities. Third, the data is often spread across multiple systems, and piecing together a causal graph is challenging. As a result these systems employ numerous heuristics and models to prevent false positives, and to make the investigative process tractable. For example, systems often prune irrelevant dependencies, and only track those that are relevant to the attack, or prioritize alerts based on the severity of the attack, and the likelihood of the alert being a true positive. Systems also use a variety of techniques to reduce the number of alerts that need to be investigated, including using graph topological properties, or rareness scores to prioritize alerts.
@@
The drawback however, is that these automated techniques assume an ideal world – one where the system has access to all the data, every event is logged, and often ignore the practical challenges in analyzing such threats (e.g., the time cycles an analyst spends on a threat). In this paper, we survey the state-of-the-art in enterprise threat response systems. We first describe the challenges in building such detection or investigation systems. We then describe the different types of research systems that have been proposed, the heuristics or models they use, and end with a discussion of open problems in this space.
@@ 2.1. Classic EDR and SIEM Systems
However, while these systems can provide good coverage in tracking low-level system events, the burden often lies on the analyst to leverage their skill, and expertise to piece together the context of an attack. Indeed, SIEM rules aid by automatically detecting signatures or rules from collected events (again, using expertly written rules). Unfortunately, though these systems form a part of the puzzle, and can often be evaded by subtly modifying the attack to avoid detection via rule-matching  [^3] Indeed Uetz et al. (Uetz et al., 2023) showed in their recent USENIX ’23 paper that nearly half of the current SIEM rules could be easily evaded by adversaries by using simple evasion tactics..
The industry’s adoption of EDR systems has been sudden and quick, with nearly 56% of organizations in 2021 having begun to deploy these within their enterprise, according to a survey (Business, 2024). As a result, enterprises today have contractual obligations with their EDR vendor (e.g., Qualys, FireEye), to deploy EDR tools in a low-cost, scalable manner within their enterprise, “plugged in” and ready to capture threats. This further compounds the fact that there is ultimately a “human-in-the-loop” to make sense of the various alert, and indeed. However, nearly 70% of organizations have reported that they struggle to keep up with the volume of alerts these EDR systems generate (Kaspersky, 2022). In parallel, a myriad of automated attack investigation systems have been proposed in recent literature to aid the process of threat investigation.
@@ 2.2. Origins of Automated Investigation
2.3. Birds-Eye View of an Attack Kill Chain
In Figure 1 we present a birds-eye view of a typical APT style attack kill chain based on a real incident report at Cloudflare (Prince et al., 2024). The kill chain starts with the attacker gaining access to three systems using unrotated credentials from the okta credential breach (Zaman et al., 2023). The attackers first authenticates with the Gateway using an access token they (step 1), and proceeds to access the Atlassian Server using a service account credential(associated with SmartSheet SaaS application). The attacker then proceeds to do reconnaissance for a period of time by accessing Atlassian tools like Jira, and Confluence, and viewing support tickets and product documentation (step 2). The attacker then creates a new user and adds them to privileged groups, as a measure to ensure their persistence in the event the SmartSheet credential gets revoked (step 3). Further, to establish persistence, a Command-and-Control (C2) backdoor software (Sliver Adversary Emulation Framework) is installed and used to exfiltrate data. Lastly, the attackers accessed BitBucket Code repositories (again using compromised credentials), and download several repositories of data. This attack was detected by the Cloudflare team (nearly a month from initial compromise, by an alert) and the user accounts were deactivated, and the C2 backdoor was removed. The attackers also failed to laterally move but access a non-production console server in a different data center (based on a non-enforced ACL) but no evidence of further access was found.
2.4. System Classification
@@
Systems that perform automated attack investigation can be classified on several parameters – data source, investigation type, system type, etc. Perhaps the broadest classification we can make is based on the detection system itself, i.e., if it is online or offline system. Online systems are those that detect attacks in a real-time (e..g, NodLink (Li et al., 2023)), while offline systems work by analyzing already collected logs or events (e.g., Airtag (Ding et al., 2023), ProvDetector (Wang et al., 2020), OmegaLog (Hassan et al., 2020b)). Based on the data source used for investigation, systems can be classified as provenance (e.g., Morse (Hossain et al., 2020), NoDoze (Hassan et al., 2019), PrioTracker (Liu et al., 2018b)), or other sources (e.g., Hopper (Ho et al., 2021), Latte (Liu et al., 2018a)). Fine-grained systems exploit provenance by tracking OS-level events, while coarse-grained systems track specific components such as Active Directory log-on events.

### v2

@@ 1. Introduction
Once inside, attackers use a variety of techniques to move laterally through the network (e.g., pass-the-hash, Kerberoasting, etc), compromise accounts with high privileges and access to sensitive systems or data, and use those privileges to ferry data outside the network. Throughout this process, attackers play a carefully crafted cat-and-mouse game against threat detection systems, trying to masquerade as legitimate users within the network (e.g., by using legitimate tools, or probing access to systems infrequently). Indeed, this process can often be long, and as a result, attackers can persist inside the network for weeks, if not months before being detected. [^2] IBM reports the average time to detect and contain a data breach in 2023 being 204 days (IBM, 2023) For example, the 2020 SolarWinds hack persisted in the network for 14 months before being detected. The 2023 HHS ransomware attack persisted in the network for 9 months before being detected. The 2023 23andMe breach took place over a duration of 5 months before being detected.
@@
Enterprises too employ various systems to detect and respond to such attacks – those that monitor endpoint or host activities (e.g., Endpoint Detection and Response systems), those that monitor network traffic (e.g., Intrusion and Detection Systems, or Intrusion Prevension Systems), monitor user activity (e.g., User and Entity Behavior Analytics), aggregate and monitor system information such as logs (e.g., Security Information and Event Management). These systems are often deployed on endpoint devices (e.g., servers, laptops), inside network infrastructures (e.g., routers, switches), or in centralized hardware (e.g., Active Directory server), and are passively monitoring for threats. While a combination of these systems is deployed in modern enterprises today, a team of analysis are tasked to triage and respond to threats these systems generate. It is this step that is labor intensive and time consuming, and often has the highest risk of missing a true positive (commonly termed the needle-in-the-haystack problem). As a result, several systems have been proposed to aid the investigative process by tracking related events, piecing together a causal reasoning of those events that led to the attack (termed ”causal analysis”), and understanding the scope of the attack, all in an effort to minimize false positives, and accelerate the triage process.
@@
Unfortunately, the scale and complexity of an enterprise poses several challenges in building such systems. First, the sheer volume of data that needs to be analyzed is large. Second, the data is often noisy, and is heavily skewed towards benign activities. Third, the data is often spread across multiple systems, and piecing together a causal graph is challenging. As a result these systems employ numerous heuristics and models to prevent false positives, and make the investigative process tractable. For example, systems often prune irrelevant dependencies, and only track those that are relevant to the attack, or prioritize alerts based on the severity of the attack, and the likelihood of the alert being a true positive. Systems also use a variety of techniques to reduce the number of alerts that need to be investigated, including using graph topological properties, or rareness scores to prioritize alerts.
@@
The drawback however is that automated investigation systems assume an ideal world – one where the system has access to all the data, every event is logged, and often ignore the practical challenges in analyzing such threats (e.g., the time cycles an analyst spends on a threat, missing or restricted access to data). In this paper, we survey the state-of-the-art in enterprise threat response systems. We first describe the challenges in building such a detection or investigation system. We then describe the different types of systems proposed by prior work, the heuristics or models they use, and end with a discussion of open problems in this space.
@@ 2.1. Classic EDR and SIEM Systems
However, while these systems can provide good coverage in tracking low-level system events, the burden often lies on the analyst to leverage their skill, and expertise to piece together the context of an attack. Indeed, SIEM rules aid by automatically detecting signatures or rules from collected events. Unfortunately, these rules aren’t bullet-proof and can often be evaded by subtly modifying the attack to avoid detection via rule-matching  [^3] Indeed Uetz et al. (Uetz et al., 2023) showed in their recent USENIX ’23 paper that nearly half of the current SIEM rules could be easily evaded by adversaries by using simple evasion tactics..
The industry’s adoption of EDR systems has been sudden and quick, with nearly 56% of organizations in 2021 having begun to deploy these within their enterprise, according to a survey (Business, 2024). As a result, enterprises today have contractual obligations with their EDR vendor (e.g., Qualys, FireEye), to deploy EDR tools in a low-cost, scalable manner within their enterprise, plugged in and ready to capture threats [^4] Often motivated by indirect reasons, e.g., lowering the premiums on your cyber insurance policy.. This further compounds the fact that there is ultimately a “human-in-the-loop” to make sense of the various alert, and indeed. However, nearly 70% of organizations have reported that they struggle to keep up with the volume of alerts these EDR systems generate (Kaspersky, 2022). In parallel, a myriad of automated attack investigation systems have been proposed in recent literature to aid the process of threat investigation.
@@ 2.2. Origins of Automated Investigation
2.3. Modern Investigation Systems
@@
Systems that perform automated attack investigation vary on several parameters – data source, investigation type, system type, etc. Perhaps the broadest classification we can make is based on the detection system itself, i.e., if it is online or offline system. Online systems are those that detect attacks in a real-time (e..g, NodLink (Li et al., 2023)), while offline systems work by analyzing already collected logs or events (e.g., Airtag (Ding et al., 2023), ProvDetector (Wang et al., 2020), OmegaLog (Hassan et al., 2020b)). Based on the data source used for investigation, systems can be classified as provenance-based (e.g., Morse (Hossain et al., 2020), NoDoze (Hassan et al., 2019), PrioTracker (Liu et al., 2018b)), or other data sources (e.g., Hopper (Ho et al., 2021), Latte (Liu et al., 2018a)). Fine-grained systems exploit provenance by tracking OS-level events, while coarse-grained systems track specific components such as Active Directory log-on events.
@@
2.4. Birds-Eye View of an Attack Kill Chain
To constrast these tools and the nature of their design, we present a modern APT style attack kill chain. Figure 1 presents a birds-eye view of an attack kill chain based on a real incident report at Cloudflare (Prince et al., 2024). The kill chain starts with the attacker gaining access to three systems using unrotated credentials from the okta credential breach (Zaman et al., 2023). The attackers first authenticates with the Gateway using an access token they (step 1), and proceeds to access the Atlassian Server using a service account credential(associated with SmartSheet SaaS application). The attacker then proceeds to do reconnaissance for a period of time by accessing Atlassian tools like Jira, and Confluence, and viewing support tickets and product documentation (step 2). The attacker then creates a new user and adds them to privileged groups, as a measure to ensure their persistence in the event the SmartSheet credential gets revoked (step 3). Further, to establish persistence, a Command-and-Control (C2) backdoor software (Sliver Adversary Emulation Framework) is installed and used to exfiltrate data. Lastly, the attackers accessed BitBucket Code repositories (again using compromised credentials), and download several repositories of data. This attack was detected by the Cloudflare team (nearly a month from initial compromise, by an alert) and the user accounts were deactivated, and the C2 backdoor was removed. The attackers also failed to laterally move but access a non-production console server in a different data center (based on a non-enforced ACL) but no evidence of further access was found.